OpenStudy (anonymous):

What is the solution of |x – 2| > –3? Explain.

OpenStudy (lgbasallote):

simply solve for x in x - 2 > - 3 and x - 2 < 3 got it?

OpenStudy (jiteshmeghwal9):

as we don't know that the value of x is negative or positive so the ans will come both in positive & negative:)

OpenStudy (dls):

x=3 because that is the minimum value to avoid negative sign in mod function basic !

OpenStudy (jiteshmeghwal9):

x=-3+2=-1 -x=-3+2=-1 x=-(-1)=1

OpenStudy (anonymous):

When ever you have an inequality like this it's always best to rewrite it as such: \[-3 < x-2 < 3\] Then just solve for x by subtracting two from everywhere in the inequality. Note: when multiplying by a negative, it will switch the inequality signs.

OpenStudy (anonymous):

Thank you guys!

OpenStudy (dls):

Please close the question :)